OpenStudy (anonymous):

Solveand check 2/1-x=3/1+x

jimthompson5910 (jim_thompson5910):

the problem is \[\Large \frac{2}{1-x} = \frac{3}{1+x}\] right?

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

\[\Large \frac{2}{1-x} = \frac{3}{1+x}\] \[\Large 2(1+x) = 3(1-x)\] \[\Large 2+2x = 3-3x\] I'll let you finish. Don't forget to check your answer.

OpenStudy (anonymous):

let me finish what do i do

jimthompson5910 (jim_thompson5910):

you solve for x by getting it to one side and everything else to the other side

jimthompson5910 (jim_thompson5910):

so you want something like x = ....

OpenStudy (anonymous):

ok I undestand that but how do i do that

jimthompson5910 (jim_thompson5910):

first move the -3x over by adding 3x to both sides

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

\[\Large 2+2x = 3-3x\] \[\Large 2+2x+3x = 3-3x+3x\] \[\Large 2+5x = 3\]

jimthompson5910 (jim_thompson5910):

now you need to move that '2' on the left side over to the right

OpenStudy (anonymous):

2+3

jimthompson5910 (jim_thompson5910):

no, you need to undo the +2 on the left side by subtracting 2 from both sides

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

thats on the left, what about the right?

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

so 5x = 1 which means x = ???

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

5 times 1 = 1 ???

OpenStudy (anonymous):

=5

jimthompson5910 (jim_thompson5910):

5 times ____ = 1 what goes in the blank

OpenStudy (anonymous):

i do not know t hrought it was 5

OpenStudy (anonymous):

4 maybe

jimthompson5910 (jim_thompson5910):

how do you undo multiplication?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

5-1=4

jimthompson5910 (jim_thompson5910):

subtraction is the opposite of addition, not multiplication

OpenStudy (anonymous):

division

jimthompson5910 (jim_thompson5910):

so you have to divide both sides by ____

OpenStudy (anonymous):

divide by 1

jimthompson5910 (jim_thompson5910):

no, you want to move that 5 over

OpenStudy (anonymous):

ok I m total confused

jimthompson5910 (jim_thompson5910):

5x = 1 you want to move the 5 over

OpenStudy (anonymous):

so move it to =5

jimthompson5910 (jim_thompson5910):

no you have to divide both sides by 5

OpenStudy (anonymous):

where what sdes

OpenStudy (anonymous):

would it be 5 divide by2 ad 5 divide by 3

jimthompson5910 (jim_thompson5910):

no just divide both sides by 5

OpenStudy (anonymous):

what side what o i divide

jimthompson5910 (jim_thompson5910):

both sides because you have to balance things out

jimthompson5910 (jim_thompson5910):

5x divided by 5 becomes what?

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

just x

jimthompson5910 (jim_thompson5910):

5x divided by 5 becomes x

OpenStudy (anonymous):

x/1

jimthompson5910 (jim_thompson5910):

yeah which is just x

OpenStudy (anonymous):

so it is x

jimthompson5910 (jim_thompson5910):

yes the left side becomes just x

jimthompson5910 (jim_thompson5910):

you must do the same to the right side

OpenStudy (anonymous):

what is the right side

jimthompson5910 (jim_thompson5910):

5x = 1 x = ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

check out this page http://www.purplemath.com/modules/solvelin2.htm

OpenStudy (anonymous):

x=5/2

jimthompson5910 (jim_thompson5910):

that's the example on the page, not the answer to this question

OpenStudy (anonymous):

so how do i get my answer

jimthompson5910 (jim_thompson5910):

here's another page that might help http://www.regentsprep.org/Regents/math/ALGEBRA/AE2/LSolvEq.htm

jimthompson5910 (jim_thompson5910):

these pages explain how to solve linear equations

OpenStudy (anonymous):

I looked at it and still not sure

jimthompson5910 (jim_thompson5910):

read it over a few times and try some of the examples on your own so you can get practice

OpenStudy (anonymous):

I do not have a lot of timebut will try

jimthompson5910 (jim_thompson5910):

well you'll have to make time somehow

OpenStudy (anonymous):

whateverbye

jimthompson5910 (jim_thompson5910):

this is very important because you'll use it later down the road and there's no need to be rude, I'm just trying to help

jimthompson5910 (jim_thompson5910):

this is NOT an answer service, we don't do your homework for you

OpenStudy (anonymous):

but you are not helping

OpenStudy (anonymous):

I do not want you to do it for me ust help me

jimthompson5910 (jim_thompson5910):

I've done enough that hopefully you can see what to do next, I don't want to do it all for you

OpenStudy (anonymous):

I am not askig you to do it

jimthompson5910 (jim_thompson5910):

you see how to go from 2x = 5 to x = 5/2 right? you divide both sides by 2

jimthompson5910 (jim_thompson5910):

you're doing the same thing with 5x = 1

OpenStudy (anonymous):

0.4,2.5

jimthompson5910 (jim_thompson5910):

leave it as a fraction

jimthompson5910 (jim_thompson5910):

5/2 is just fine the way it is but focus on 5x = 1

OpenStudy (anonymous):

0.4/2.5

jimthompson5910 (jim_thompson5910):

no i don't know how you're getting that

OpenStudy (anonymous):

divided 2from 5

jimthompson5910 (jim_thompson5910):

ok don't worry about converting 5/2 into a decimal

OpenStudy (anonymous):

is it 5/2

jimthompson5910 (jim_thompson5910):

5/2 is the answer to 2x = 5 but not 5x = 1

OpenStudy (anonymous):

would i be 5

jimthompson5910 (jim_thompson5910):

no replace x with 5 5x = 1 5 times 5 = 1 .... is that true??

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

so the answer isn't x = 5

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

5x = 1 5 times 4 = 1 .... that's not true either

OpenStudy (anonymous):

5 divided 5 is 1

jimthompson5910 (jim_thompson5910):

good, so that's the left side what about the right?

OpenStudy (anonymous):

what is the right side

jimthompson5910 (jim_thompson5910):

you divided the left side by 5 you have to do the same to the right side

OpenStudy (anonymous):

5 divided by what

jimthompson5910 (jim_thompson5910):

you divided the left side 5x by 5 to get x you have to do the same to the right side

jimthompson5910 (jim_thompson5910):

so the right side of 1 becomes _____

OpenStudy (anonymous):

5

OpenStudy (anonymous):

3 divided by 5

jimthompson5910 (jim_thompson5910):

where are you getting 3?

OpenStudy (anonymous):

on the right side

jimthompson5910 (jim_thompson5910):

it's 5x = 1, not 5x = 3